OpenStudy (anonymous):

Given the equation 6X-y=98, find x if y = –2. A. 16 B. 100/6 C. 86 D. 110

OpenStudy (anonymous):

put y=-2 in the given equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u will get 6x+2=98

OpenStudy (anonymous):

now 6x+2-2=98-2 subtract 2 from both side

OpenStudy (anonymous):

u will get 6x=96

OpenStudy (anonymous):

x=96/6

OpenStudy (anonymous):

alright so it will be A

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cuz i divide 96 and 6 it give me 16

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yes u are right